# XUE: The CO2-rich terrestrial planet-forming region of an externally irradiated Herbig disk

### v1

@@ 5.2.2 Transport-dependent chemistry
We start by considering the process of gas advection, connected to the dust drift, which leads to viscous gas accretion onto the star. While the m-size dust grains are well coupled to the gas and follow its motion, the bigger grains, so-called pebbles, will decouple from the gas and radially drift inward, settling towards the disk midplane. Due to the gradient in temperature across the disk, and accounting for some vertical mixing, these pebbles will release part of their icy mantle in the gas-phase via sublimation when crossing the ice line of their mantle components (Ciesla & Cuzzi, 2006; Lambrechts & Johansen, 2012; Banzatti et al., 2020, 2023). The main ice lines in protoplanetary disks correspond, in order of increasing radial distance from the star, to H2O, CO2, CO, and methane (CH4), but their exact radial and vertical positions are determined by the detailed thermal structure of the disk (Öberg et al., 2011; Piso et al., 2015).
@@
We may therefore be observing the CO2-rich gas phase of XUE 10 that follows the advection of the bulk of water vapor and its viscous accretion onto the star, i.e., while the CO2 gas is still drifting inward. This may be reasonable, considering that the advection timescale is expected to be shorter in IMTT/Herbig disks compared to lower-mass counterparts, as the accretion rates are generally higher, and the gas temperatures may enhance the disk viscosity, and therefore the inward advection (e.g., Dullemond & Dominik, 2004; Calvet et al., 2004; Garcia Lopez et al., 2006). However, the VLT/KMOS infrared spectrum of XUE 10 (Ramírez-Tannus et al., 2020) does not show evidence of the Br line, one of the main proxies for stellar accretion, suggesting that XUE 10 has a low accretion luminosity, and therefore gas advection may not be very efficient. There is also evidence that midplane cosmic-ray UV fields can boost OH and favor CO2‑rich ice formation in the inner disk. As these pebbles drift inward and their volatiles desorb, the resulting gas can absorb 30% of the cosmic-ray UV flux, throttling new H2O photo-dissociation and locking in the CO2 excess (Molano & Kamp, 2012a, b).
@@
Even if the water vapor was being effectively advected past the snow line, we might be insensitive to an enhancement in column density of gas-phase H2O in the upper disk layers with infrared observations. In fact, the influx of dust that accompanies the inward drift of water-ice may obscure the enhanced emission of water vapor near the snow line (Sellek et al., 2025; Houge et al., 2025). This is because the dry dust grains inside the snow line are thought to fragment more easily; the resulting smaller grains will couple to the gas, producing a ?traffic jam? of dust (Pinilla et al., 2016), and increase the dust opacity () in the inner disk. Outside the H2O snow line, instead, the icy dust grains are predicted to have a higher fragmentation velocity, i.e., they fragment less easily (Blum & Wurm, 2008; Gundlach & Blum, 2015). Therefore, the CO2 emission coming from the ice line further out in the disk will be less affected by this increase of dust opacity.
@@
Following the discussion in Bosman et al. (2017), we highlight that the line ratio between the 12CO2 and 13CO2 isotopic species is informative of the ice sublimation at the CO2 ice line. As the -branch of 13CO2 visible at 15.42 m is optically thinner than 12CO2, it may trace deeper in the disk, and therefore be more sensitive to the abundance structure of CO2. We note that a similar argument applies to other optically thinner isotopic species, i.e., 16O12C18O and 16O12C17O in this context. At present, however, these isotopic species are more challenging to compare using line ratios, given their unique simultaneous detections in XUE 10, and their expected much lower abundance and band strength. Moreover, it is unclear whether these isotopologues are tracing deeper disk layers than 12CO2 (see Section 4.2).
@@
Focusing on 13CO2, its -branch largely dominates over the nearby contributions from the CO2 -branch in the spectral region of XUE 10 of Figure 8, with a peak line flux of 18 mJy. Similarly to Grant et al. (2023), in XUE 10 the CO2 (25) -branch line (15 mJy) at 15.45 m is stronger than the (23) line (12 mJy) at 15.41 m, due to the contribution of a hot band of the -branch that arises only at high temperatures or high column densities. We therefore considered the (27) line (7 mJy) at 15.48 m as representative of individual -branch lines with similar levels. From our best-fit models, we estimate a peak line ratio between the 13CO2 -branch and the (27) line of 12CO2 to be 2.37 1.05, which is comparable within error bar to the ratio found in GW Lup (1.4; Grant et al., 2023), and MY Lup (2.24 0.24, Salyk et al., 2025). The non-LTE CO2 excitation model computed with DALI by Bosman et al. (2017) predicts that a different initial inner/outer abundance (/) of CO2 (with respect to the total gas density), or a different gas-to-dust ratio () can impact the observed line ratios between 12CO2 - and - lines, and between the 13CO2 -branch and a representative 12CO2 - line. The -branch/(27) line ratio we measure towards XUE 10 is consistent within uncertainties with a wide range of models (see Figure 9 of Bosman et al., 2017). This includes the case of GW Lup, which points at a low outer ( = 10-8) and high inner ( = 10-6) disk CO2 abundance at enhanced gas-to-dust ratio ( = 10 000). This finding suggests that the difference in optical depth between 12CO2 and 13CO2 measured towards XUE 10 is not enough to break the degeneracy between different DALI models. A non-LTE treatment of these molecular species in the mid-IR spectrum of XUE 10 may also be required.
@@
In this second scenario, we assumed so far a smooth transport of gas and dust through the inner disk, where the H2O vapor is removed by advection and by stellar UV photo-dissociation, while CO2 vapor remains abundant in the gas-phase and can be strongly enhanced by local gas-phase production from OH and CO in cold gas. Nevertheless, there is a number of processes that can slow down or impede this inward transport of gas and dust, and they may play a relevant role in this scenario. We will focus on the presence of a dust trap or a gas cavity in the inner/outer disk, and outer disk truncation by photoevaporation.
@@
The presence of a inner dust trap or a gas cavity between the H2O and the CO2 ice lines has already been proposed to explain a gas-phase H2O–CO2 dichotomy in the inner disk (Gasman et al., 2023; Grant et al., 2023; Vlasblom et al., 2025). While the H2O vapor will be suppressed across the gas cavity, or while the water ice-rich grains will be trapped beyond the snow line in the dust trap, CO2 will still be abundant in the lukewarm gas phase, and its line flux will be less affected, because the contributions to its emission come from larger disk radii (Walsh et al., 2015; Anderson et al., 2021; Sellek et al., 2025). Dust traps and gas cavities may also be present in the outer disk (if not truncated), and similarly act against the delivery of gas and water ice-rich dust grains to the inner disk. Other gap-opening mechanisms may be the internal X-ray photoevaporation driven by the central star, and forming planets in the disk (Lienert et al., 2024). While we cannot draw any conclusion regarding the latter, internal X-ray photoevaporation is unlikely to play a role here, given that XUE 10 is undetected in X-ray with Chandra (X 1029 erg s-1).
@@
The spectral energy distribution of XUE 10 is consistent with Group I Herbig ?flaring? disks (Ramírez-Tannus et al., 2025), which are often characterized by large cavities (e.g., Banzatti et al., 2018), favoring the hypothesis of the presence of substructures in this disk. However, the location of traps/cavities must be compared to the location of the ice lines. Due to the higher stellar luminosity of IMTT and Herbig stars compared to T Tauris, the ice lines of H2O and CO2 will be pushed significantly outward in a disk like XUE 10. In a flaring disk passively irradiated by a central star with the luminosity of XUE 10 ( 69 ⊙), we estimate that the H2O and CO2 ice lines should be roughly located at 21 au and 85 au in radius respectively (Chiang & Goldreich, 1997). For comparison, Grant et al. (2023) estimate in GW Lup ice line locations at 0.4 au (H2O) and at a slightly larger radius than 1 au (CO2). Therefore, the trap/cavity hypothesis may explain the CO2–rich spectrum of XUE 10, but requiring that the outer disk is not truncated below 100 au, which may actually be the case for the XUE sample. High angular resolution observations are needed to spatially resolve the inner and outer disk and look for substructures, which are common among nearby Herbig disks (e.g., Andrews et al., 2018). This type of observation is almost unfeasible towards XUE 10 with the current generation of telescopes, but it will become possible in the near future with the Extremely Large Telescope (ELT)444https://elt.eso.org, and with the ALMA Wideband Sensitivity Upgrade (WSU)555https://www.almaobservatory.org/en/scientists/alma-2030-wsu/wsu-program/.
@@
If the outer disk was to be truncated by the external FUV irradiation, similar to what is observed in disks in the ONC (e.g., Boyden & Eisner, 2020, 2023), this could result in gas depletion (Keyte & Haworth, 2025), and also in a reduced influx of dust into the inner disk, cutting down the reservoir of drifting ice-rich dust grains (Facchini et al., 2016; Qiao et al., 2023; Gárate et al., 2024; Paine et al., 2025). Although the influx of CO2 ice (and gas) would be reduced in a similar way to water, the combination of an outer drainage of gas and dust with the in-situ chemistry discussed in Section 5.2.1 may reproduce the observed CO2-rich spectrum of XUE 10. The recently developed model by Sellek et al. (2025), which considers the 1D inner disk chemical evolution resulting from gas viscous evolution and dust radial drift, also in presence of traps/cavities, represents an optimal way to further benchmark the transport of H2O and CO2 in IMTT/Herbig disks similar to XUE 10. Hereby, the difference in advection timescale, dust opacity, and depletion mechanism of H2O compared to T Tauri sources may crucially impact on the H2O and CO2 observables in the inner regions of IMTT and Herbig disks, and validate one of our proposed explanations.
@@ 5.3 Oxygen isotopic ratios in relation with Solar System meteorites and the ISM
In Section 4.2 we concluded that 13CO2 is most likely emitting at 12C/13C ISM ratio. Hence, in the following we will focus on the interpretation of potential oxygen isotopic anomalies detected in 16O12C18O and 16O12C17O in a protoplanetary disk.
@@
An extensive literature over five decades discusses oxygen isotopic anomalies in meteorites, i.e., solids that condensed in the Sun’s protoplanetary disk about 4.53 Gyr ago (Clayton, 1993). Meteoritic solids are thought to have interacted with gaseous reservoirs in the Solar Nebula, when the protosolar disk was still present. The isotopic anomalies found in meteorites are often too large to be explained by atomic mass-dependent fractionation, i.e., according to the mass of the isotope, or aqueous processing in the disk. A mass-independent isotopic fractionation in the gas reservoir is required. In particular, the reservoir of oxygen isotopes in the Solar Nebula appears to have experienced complex spatial and temporal variations, as observed in various classes of carbonaceous chondrites, i.e., chondrule-poor stony meteorites (Taylor et al., 1965; Clayton et al., 1973; Clayton & Mayeda, 1999), and early and late crystallizing minerals (Clayton, 1993). Also the mobility of meteoritic components within the disk may have played a role in the Solar Nebula, e.g., they might have been ejected from inner to outer regions in an X-ray wind, as discussed by Shu et al. (2000).
@@
Various processes have been proposed to coherently explain the meteoritic findings, e.g., isotopic inheritance from different parts of the interstellar medium, supernova ejecta, spallation by solar flare protons in the disk, and disequilibrium chemistry. Thiemens & Heidenreich (1983) were the first to make a connection between a source of UV radiation, particularly the young Sun, with the mass-independent fractionation of 18O/17O oxygen atoms in the O2 molecule. Later on, Lyons & Young (2005) proposed the selective photo-dissociation of CO, rather than O2, caused by the UV irradiation of a young O or B star external to the Solar Nebula to explain the mass-independent fractionation of oxygen atoms in meteorites. The scenario of CO photo-dissociation caused by the UV irradiation of the central star has been validated for protoplanetary disks by Visser et al. (2009) and Miotello et al. (2014).
@@
Specifically, in the outer regions of gas-rich disks, at 1015 cm-2 CO can self-shield against UV irradiation, leading to a depletion of 12C18O and 12C17O relative to 12C16O in the gas-phase. This is because the UV radiation penetrates deeper in the optically thinner absorbing lines of the less abundant isotopologues. The relatively high 12C16O/12C18O ratio observed in CO towards a small sample of nearby T Tauri disks (Brittain et al., 2005; Smith et al., 2009) may be explained by this isotopic-selective CO self-shielding. In XUE 10, though, we find enhanced chemical abundances of the 16O12C18O and 16O12C17O isotopologues of CO2 instead of CO. A transport-dependent chemical scenario can account for the observed oxygen isotopic anomaly in CO2 in the inner disk of a protoplanetary disk. Assuming that CO2 ices have a standard (ISM) isotopic composition in oxygen, the isotopic anomaly can be transferred to CO2 from CO gas and H2O ices via photochemistry. In particular, isotope-selective UV photo-dissociation of CO in the outer disk releases 18O and 17O atoms that can be vertically mixed downward in the disk through the ?cold finger? effect. These oxygen atoms can then participate in surface reactions on dust grains in the midplane, forming HO- and HO-enriched water ices (Stevenson & Lunine, 1988; Meijerink et al., 2009). These isotopically enriched water ices then need to be transported inward, possibly via radial drift of dust grains. Near the snow line, accounting for vertical mixing, the ices sublimate and water photo-dissociation can release 18OH and 17OH atoms into the gas-phase, allowing in the end the formation of large amounts of 16O12C18O and 16O12C17O vapor.
@@
Alternatively, CO2 ices may be intrinsically anomalous in their oxygen composition, not reflecting an anomaly in H2O ice. Accounting for inward dust drift, these CO2 ices will eventually cross the CO2 ice line, releasing 16O12C18O and 16O12C17O into the gas phase and enriching the inner disk reservoir of O isotopes locked in CO2. Yet, the abundance of icy CO2 transported inward is expected to be significantly lower than that of H2O ice. Consequently, the enrichment of the inner disk gas reservoir following ice line sublimation is likely dominated by water, unless it has already been accreted onto the central star (Sellek et al., 2025). Nevertheless, an anomalous strength of infrared molecular bands may not reflect anomalies in the abundance structure of a given molecular species in the disk, as different spectral features may emit from different disk layers. A meaningful translation between CO2 infrared band strengths and abundance pattern requires a detailed isotope-sensitive thermochemical-structural model for this source.

### v2

@@ 5.2.2 Transport-dependent chemistry
We start by considering the process of gas advection, connected to the dust drift, which leads to viscous gas accretion onto the star. While the m-size dust grains are well coupled to the gas and follow its motion, the bigger grains, so-called pebbles, will decouple from the gas and radially drift inward, settling toward the disk midplane. Due to the gradient in temperature across the disk, and accounting for some vertical mixing, these pebbles will release part of their icy mantle in the gas phase via sublimation when crossing the ice line of their mantle components (Ciesla & Cuzzi, 2006; Lambrechts & Johansen, 2012; Banzatti et al., 2020, 2023). The main ice lines in protoplanetary disks correspond, in order of increasing radial distance from the star, to H2O, CO2, CO, and methane (CH4), but their exact radial and vertical positions are determined by the detailed thermal structure of the disk (Öberg et al., 2011; Piso et al., 2015).
@@
We may therefore be observing the CO2-rich gas phase of XUE 10 that follows the advection of the bulk of water vapor and its viscous accretion onto the star, i.e., while the CO2 gas is still drifting inward. This may be reasonable, considering that the advection timescale is expected to be shorter in IMTT or Herbig disks compared to lower-mass counterparts, as the accretion rates are generally higher, and the gas temperatures may enhance the disk viscosity and therefore the inward advection (e.g., Dullemond & Dominik, 2004; Calvet et al., 2004; Garcia Lopez et al., 2006). However, the VLT/KMOS infrared spectrum of XUE 10 (Ramírez-Tannus et al., 2020) does not show evidence of the Br line, one of the main proxies for stellar accretion, suggesting that XUE 10 has a low accretion luminosity, and therefore gas advection may not be very efficient. There is also evidence that midplane cosmic-ray UV fields can boost OH and favor CO2-rich ice formation in the inner disk. As these pebbles drift inward and their volatiles desorb, the gas can absorb 30% of the cosmic-ray UV flux, throttling new H2O photo-dissociation and locking in the CO2 excess (Molano & Kamp, 2012a, b).
@@
Even if the water vapor was being effectively advected past the snow line, we might be insensitive to an enhancement in column density of gas phase H2O in the upper disk layers with infrared observations. In fact, the influx of dust that accompanies the inward drift of water-ice may obscure the enhanced emission of water vapor near the snow line (Sellek et al., 2025; Houge et al., 2025). This is because the dry dust grains inside the snow line are thought to fragment more easily; the resulting smaller grains will couple to the gas, producing a ?traffic jam? of dust (Pinilla et al., 2016), and increase the dust opacity () in the inner disk. Outside the H2O snow line, instead, the icy dust grains are predicted to have a higher fragmentation velocity, i.e., they fragment less easily (Blum & Wurm, 2008; Gundlach & Blum, 2015). Therefore, the CO2 emission coming from the ice line further out in the disk will be less affected by this increase of dust opacity.
@@
Following the discussion in Bosman et al. (2017), we highlight that the line ratio between the 12CO2 and 13CO2 isotopic species is informative of the ice sublimation at the CO2 ice line. As the -branch of 13CO2 visible at 15.42 m is optically thinner than 12CO2, it may trace deeper in the disk and therefore be more sensitive to the abundance structure of CO2. We note that a similar argument applies to other optically thinner isotopic species, i.e., 16O12C18O and 16O12C17O in this context. At present, however, these isotopic species are more challenging to compare using line ratios, given their unique simultaneous detections in XUE 10, and their expected much lower abundance and band strength. Moreover, it is unclear whether these isotopologues are tracing deeper disk layers than 12CO2 (see Section 4.2).
@@
Focusing on 13CO2, its -branch largely dominates over the nearby contributions from the CO2 -branch in the spectral region of XUE 10 of Figure 8, with a peak line flux of 18 mJy. Similarly to Grant et al. (2023), in XUE 10 the CO2 (25) -branch line (15 mJy) at 15.45 m is stronger than the (23) line (12 mJy) at 15.41 m, due to the contribution of a hot band of the -branch that arises only at high temperatures or high column densities. We therefore considered the (27) line (7 mJy) at 15.48 m as representative of individual -branch lines with similar levels. From our best-fit models, we estimate a peak line ratio between the 13CO2 -branch and the (27) line of 12CO2 to be 2.37 1.05, which is comparable within error bar to the ratio found in GW Lup (1.4; Grant et al., 2023), and MY Lup (2.24 0.24, Salyk et al., 2025). The non-LTE CO2 excitation model computed with DALI by Bosman et al. (2017) predicts that a different initial inner/outer abundance (/) of CO2 (with respect to the total gas density), or a different gas-to-dust ratio () can impact the observed line ratios between 12CO2 - and - lines, and between the 13CO2 -branch and a representative 12CO2 - line. The -branch/(27) line ratio we measure toward XUE 10 is consistent within uncertainties with a wide range of models (see Figure 9 of Bosman et al., 2017). This includes the case of GW Lup, which points at a low outer ( = 10-8) and high inner ( = 10-6) disk CO2 abundance at enhanced gas-to-dust ratio ( = 10 000). This finding suggests that the difference in optical depth between 12CO2 and 13CO2 measured toward XUE 10 is not enough to break the degeneracy between different DALI models. A non-LTE treatment of these molecular species in the mid-IR spectrum of XUE 10 may also be required.
@@
In this second scenario, we assumed so far a smooth transport of gas and dust through the inner disk, where the H2O vapor is removed by advection and by stellar UV photo-dissociation, while CO2 vapor remains abundant in the gas phase and can be strongly enhanced by local gas phase production from OH and CO in cold gas. Nevertheless, there is a number of processes that can slow down or impede this inward transport of gas and dust, and they may play a relevant role in this scenario. We focus on the presence of a dust trap or a gas cavity in the inner or outer disk, and outer disk truncation by photoevaporation.
@@
The presence of a inner dust trap or a gas cavity between the H2O and the CO2 ice lines has already been proposed to explain a gas phase H2O–CO2 dichotomy in the inner disk (Gasman et al., 2023; Grant et al., 2023; Vlasblom et al., 2025). While the H2O vapor will be suppressed across the gas cavity, or while the water ice-rich grains will be trapped beyond the snow line in the dust trap, CO2 will still be abundant in the lukewarm gas phase, and its line flux will be less affected, because the contributions to its emission come from larger disk radii (Walsh et al., 2015; Anderson et al., 2021; Sellek et al., 2025). Dust traps and gas cavities may also be present in the outer disk (if not truncated), and similarly act against the delivery of gas and water ice-rich dust grains to the inner disk. Other gap-opening mechanisms may be the internal X-ray photoevaporation driven by the central star, and forming planets in the disk (Lienert et al., 2024). While we cannot draw any conclusion regarding the latter, internal X-ray photoevaporation is unlikely to play a role here, given that XUE 10 is undetected in X-ray with Chandra (X 1029 erg s-1).
@@
The spectral energy distribution of XUE 10 is consistent with Group I Herbig ?flaring? disks (Ramírez-Tannus et al., 2025), which are often characterized by large cavities (e.g., Banzatti et al., 2018), favoring the hypothesis of the presence of substructures in this disk. However, the location of traps/cavities must be compared to the location of the ice lines. Due to the higher stellar luminosity of IMTT and Herbig stars compared to T Tauris, the ice lines of H2O and CO2 will be pushed significantly outward in a disk similar to XUE 10. In a flaring disk passively irradiated by a central star with the luminosity of XUE 10 ( 69 ⊙), we estimate that the H2O and CO2 ice lines should be roughly located at 21 au and 85 au in radius respectively (Chiang & Goldreich, 1997). For comparison, Grant et al. (2023) estimate in GW Lup ice line locations at 0.4 au (H2O) and at a slightly larger radius than 1 au (CO2). Therefore, the trap/cavity hypothesis may explain the CO2–rich spectrum of XUE 10, but requiring that the outer disk is not truncated below 100 au, which may actually be the case for the XUE sample. High angular resolution observations are needed to spatially resolve the inner and outer disk and look for substructures, which are common among nearby Herbig disks (e.g., Andrews et al., 2018). This type of observation is almost unfeasible toward XUE 10 with the current generation of telescopes, but it will become possible in the near future with the Extremely Large Telescope (ELT),444https://elt.eso.org and with the ALMA Wideband Sensitivity Upgrade (WSU).555https://www.almaobservatory.org/en/scientists/alma-2030-wsu/wsu-program/
@@
If the outer disk was to be truncated by the external FUV irradiation, similar to what is observed in disks in the ONC (e.g., Boyden & Eisner, 2020, 2023), this could result in gas depletion (Keyte & Haworth, 2025), and also in a reduced influx of dust into the inner disk, cutting down the reservoir of drifting ice-rich dust grains (Facchini et al., 2016; Qiao et al., 2023; Gárate et al., 2024; Paine et al., 2025). Although the influx of CO2 ice (and gas) would be reduced in a similar way to water, the combination of an outer drainage of gas and dust with the in situ chemistry discussed in Section 5.2.1 may reproduce the observed CO2-rich spectrum of XUE 10. The recently developed model by Sellek et al. (2025), which considers the 1D inner disk chemical evolution resulting from gas viscous evolution and dust radial drift, also in presence of traps/cavities, represents an optimal way to further benchmark the transport of H2O and CO2 in IMTT or Herbig disks similar to XUE 10. Hereby, the difference in advection timescale, dust opacity, and depletion mechanism of H2O compared to T Tauri sources may crucially impact on the H2O and CO2 observables in the inner regions of IMTT and Herbig disks, and validate one of our proposed explanations.
@@ 5.3 Oxygen isotopic ratios in relation with Solar System meteorites and the ISM
In Section 4.2 we concluded that 13CO2 is most likely emitting at 12C/13C ISM ratio. Hence, in the following we focus on the interpretation of potential oxygen isotopic anomalies detected in 16O12C18O and 16O12C17O in a protoplanetary disk.
@@
The extensive literature of over five decades discusses oxygen isotopic anomalies in meteorites, i.e., solids that condensed in the Sun’s protoplanetary disk about 4.53 Gyr ago (Clayton, 1993). Meteoritic solids are thought to have interacted with gaseous reservoirs in the solar nebula, when the protosolar disk was still present. The isotopic anomalies found in meteorites are often too large to be explained by atomic mass-dependent fractionation, i.e., according to the mass of the isotope, or aqueous processing in the disk. A mass-independent isotopic fractionation in the gas reservoir is required. In particular, the reservoir of oxygen isotopes in the solar nebula appears to have experienced complex spatial and temporal variations, as observed in various classes of carbonaceous chondrites, i.e., chondrule-poor stony meteorites (Taylor et al., 1965; Clayton et al., 1973; Clayton & Mayeda, 1999), and early and late crystallizing minerals (Clayton, 1993). The mobility of meteoritic components within the disk may have also played a role in the solar nebula; for example, meteoritic components have been ejected from inner to outer regions in an X-ray wind, as discussed by Shu et al. (2000).
@@
Various processes have been proposed to coherently explain the meteoritic findings, e.g., isotopic inheritance from different parts of the ISM, supernova ejecta, spallation by solar flare protons in the disk, and disequilibrium chemistry. Thiemens & Heidenreich (1983) were the first to make a connection between a source of UV radiation, particularly the young Sun, with the mass-independent fractionation of 18O/17O oxygen atoms in the O2 molecule. Later on, Lyons & Young (2005) proposed the selective photo-dissociation of CO, rather than O2, caused by the UV irradiation of a young O or B star external to the solar nebula to explain the mass-independent fractionation of oxygen atoms in meteorites. The scenario of CO photo-dissociation due to the UV irradiation of the central star has been validated for protoplanetary disks by Visser et al. (2009) and Miotello et al. (2014).
@@
Specifically, in the outer regions of gas-rich disks, at 1015 cm-2 CO can self-shield against UV irradiation, leading to a depletion of 12C18O and 12C17O relative to 12C16O in the gas phase. This is because the UV radiation penetrates deeper in the optically thinner absorbing lines of the less abundant isotopologues. The relatively high 12C16O/12C18O ratio observed in CO toward a small sample of nearby T Tauri disks (Brittain et al., 2005; Smith et al., 2009) may be explained by this isotopic-selective CO self-shielding. In XUE 10, though, we find enhanced chemical abundances of the 16O12C18O and 16O12C17O isotopologues of CO2 instead of CO. A transport-dependent chemical scenario can account for the observed oxygen isotopic anomaly in CO2 in the inner disk of a protoplanetary disk. Assuming that CO2 ices have a standard (ISM) isotopic composition in oxygen, the isotopic anomaly can be transferred to CO2 from CO gas and H2O ices via photochemistry. In particular, isotope-selective UV photo-dissociation of CO in the outer disk releases 18O and 17O atoms that can be vertically mixed downward in the disk through the ?cold finger? effect. These oxygen atoms can then participate in surface reactions on dust grains in the midplane, forming HO- and HO-enriched water ices (Stevenson & Lunine, 1988; Meijerink et al., 2009). These isotopically enriched water ices then need to be transported inward, possibly via radial drift of dust grains. Near the snow line, accounting for vertical mixing, the ices sublimate and water photo-dissociation can release 18OH and 17OH atoms into the gas phase, allowing in the end for the formation of large amounts of 16O12C18O and 16O12C17O vapor.
@@
Alternatively, CO2 ices may be intrinsically anomalous in their oxygen composition, and thus may not reflect an anomaly in H2O ice. Accounting for inward dust drift, these CO2 ices will eventually cross the CO2 ice line, releasing 16O12C18O and 16O12C17O into the gas phase and enriching the inner disk reservoir of O isotopes locked in CO2. Yet, the abundance of icy CO2 transported inward is expected to be significantly lower than that of H2O ice. Consequently, the enrichment of the inner disk gas reservoir following ice line sublimation is likely dominated by water, unless it has already been accreted onto the central star (Sellek et al., 2025). Nevertheless, an anomalous strength of infrared molecular bands may not reflect anomalies in the abundance structure of a given molecular species in the disk, as different spectral features may emit from different disk layers. A meaningful translation between CO2 infrared band strengths and abundance pattern requires a detailed isotope-sensitive thermochemical-structural model for this source.